OpenStudy (anonymous):

having trouble trying to integrate sqrt (900-300(t)sqrt 2+100t^2) from 0 to 3sqrt 2

myininaya (myininaya):

does that sqrt go over the whole thing?

OpenStudy (anonymous):

ya.

myininaya (myininaya):

k give me a sec let me think

myininaya (myininaya):

do you know trig substitution?

OpenStudy (anonymous):

ya but was having trouble tring to find the right thing to plug in

myininaya (myininaya):

=int(sqrt(100)*sqrt(t^2-3sqrt(2)t+9) =int(10*sqrt(t^2-3sqrt(2)t+9/2+9-9/2) =int(10*sqrt([t-3sqrt(2)/2)]^2+9/2) let tan pheta=[t-3sqrt(2)/2]/[3/sqrt(2)] {3/sqrt(2)} *tan pheta=t-3sqrt(2)/2

myininaya (myininaya):

{3/sqrt(2)}*sec^2(pheta) d(pheta)=dt so we have =int(10*(sqrt([9/2]tanpheta+9/2)*[3/sqrt(2)*sec^2pheta)

myininaya (myininaya):

10int(sqrt(9/2)*sqrt(tanpheta+1)*[3/(sqrt(2))]sec^2pheta) =10(3/sqrt(2))(3/sqrt(2))int(sqrt(sec^2pheta)sec^2pheta) =10(9/2)int(secpheta*sec^2pheta) =45int(sec^3pheta)

myininaya (myininaya):

is this helpful?

OpenStudy (anonymous):

very helpful

myininaya (myininaya):

should i go further? i havent done anything with the limits yet

OpenStudy (anonymous):

na i can do the limits

myininaya (myininaya):

k cool have fun

OpenStudy (anonymous):

thx for the help

myininaya (myininaya):

i had to complete the square above just so you know